Calculate the multifactor productivity of one of the products and services of bank , you must guess some numbers
Answer #1

Let me consider, the Insurance or Financial products associated with one output in term of USD Value and Inputs like Cost of Human Resources and Cost of Physical resources , here Multifactor productivity is calculated over 2 years - Year 2014 and Year 2015

Year
2014 2015
Output - USD 560500 678500
Number of people employed 25 30
Working hours per day 7.5 7.5
Working days per annum 240 245
Physical Resources cost 125000 175000
Average Hourly Rate for each people 30 30
Multi-Factor Productivity Formula - Output/Input 0.380 0.371

Formula

1 B C D
2 Year
3 2014 2015
4 Output - USD 560500 678500
5 Number of people employed 25 30
6 Working hours per day 7.5 7.5
7 Working days per annum 240 245
8 Physical Resources cost 125000 175000
9 Average Hourly Rate for each people 30 30
10 MultiFactor Productivity Formula - Output/Input =C4/((C8)+(C5*C6*C7*C9)) =D4/((D8)+(D5*D6*D7*D9))